# Datalog-Expressibility for Monadic and Guarded Second-Order Logic

### v2

@@ 6.3 Nested Queries
The expressive power of Datalog fragments can sometimes be increased by considering nesting, which is a concept introduced in [42] and further studied in [18]. The idea is that in nested queries we allow the use of other nested queries as if they were atomic formulas. Depending on the query language we start with, nesting may or may not increase the expressive power. Nested Datalog programs, for example, can be rewritten into equivalent Datalog programs without nesting. Similarly, MSO and GSO are closed under nesting. On the other hand, nested monadic queries (Nemodeq) are strictly more expressive than monadic queries [42], and similarly nested guarded queries (GQ+) are strictly more expressive than guarded queries, as we will see below (Proposition 6.30). We mention that the layered tree programs studied in [21] are in fact the queries obtained from linear monadic Datalog programs with at most one EDB per rule via nesting.
@@
Let $q,m\in{\mathbb{N}}$ and let $\tau$ be a finite relational signature. A $q$-nested (monadic, guarded) flag-and-check $\tau$-program of arity $m$ is defined inductively as follows. A $1$-nested (monadic, guarded) $\tau$-FCP of arity $m$ is the same as a (monadic, guarded) $\tau$-FCP of arity $m$ as defined in Definition 6.8. A $q+1$-nested (monadic, guarded) $\tau$-FCP of arity $m$ is a (monadic, guarded) $\tau$-FCP $\Pi$ of arity $m$ that may use in rule bodies $q$-nested (monadic, guarded) $\tau$-FCPs of arity $\ell$ in addition to $\ell$-ary relation symbols. These queries are called the subqueries of $\Pi$ at nesting level $q$; the subqueries of $\Pi$ at nesting level $j$ are inductively defined as the subqueries of the subqueries at nesting level $j-1$. In the case of guarded $q+1$-nested FCPs, the $q$-nested FCPs may not serve as guards. An FCP has width $m$ if all its subqueries have arity at most $m$.
@@
Note that all the classes of queries defined in Definition 6.16 are closed under disjunction by appropriately taking the union of the rule sets, just as in the proof of Lemma 4.3 for Datalog.
The class of nested guarded queries is denoted by GQ+, and the class of nested monadic queries by Nemodeq. It is known that every nested monadic query is equivalent to a Datalog program [42]; the same is even true for every nested guarded query [18, Figure 1]. The following fact is implicit in [18].
@@
  • replacing each atomic formula $R(\bar{y})$ by a new atomic formula $R^{\prime}(\bar{x}_{1},\dots,x_{m},\bar{y})$ where $R^{\prime}$ is a new IDB.
For $q>1$, the translation is defined recursively: let $\Pi_{1},\dots,\Pi_{s}$ be the subqueries of $\Pi$ at nesting level $q-1$. Let $\Pi^{\prime}_{1},\dots,\Pi^{\prime}_{s}$ be the Datalog programs that can be associated to these FCPs by the inductive assumption; we may assume that the IDBs for all these programs are pairwise disjoint. Then $\Pi^{\prime}$ is defined to be the union over the rules of $\Pi_{1},\dots,\Pi_{s}$ together with the rules of $\Pi$ where we replace each subquery $\Pi_{i}$ at nesting level $q$ by $P_{\Pi_{i}}$.
@@
Example 6.15 can be expressed in Nemodeq [42], and hence is in Datalog and in MSO.
@@ 6.4 The Nested Guarded Game
In this section we present a modification of the existential guarded $k$ pebble game that allows to capture the expressive power of GQ and GQ+. In particular, we will prove that the query in Example 6.15 is not in GQ. We call this game the nested guarded game; it will also be used in the next section to prove that there are problems in the intersection of Datalog and GSO that cannot be expressed in GQ+.
Again we start with an informal description of the game. Let $\tau$ be a finite relational signature. The game has three parameters, $q$, $m$, and $k$. There are two players, Spoiler and Duplicator, that play on a pair of $\tau$-structures $(\mathfrak{A},\mathfrak{B})$. Each player has $k$ labelled pebbles, out of which $m$ pebbles are blue, the others are red. Spoiler places the blue pebbles on $\mathfrak{A}$, Duplicator answers by placing her blue pebbles on elements of $\mathfrak{B}$. Then the two players play the existential guarded $k-m$ pebble game with the red vertices, with the difference that Duplicator looses if the map between all the pebbled vertices (blue and red alike) is not a partial homomorphism. At most $q$ times, Spoiler can relocate the red pebbles (without the guard restriction). We present a formal definition of winning strategies for Duplicator in the nested guarded game; all statements about the game only involve such winning strategies for Duplicator.
@@
Let $k,q,m\in{\mathbb{N}}$. A winning strategy for Duplicator for the $q$ nested width $m$ guarded $k$ pebble game on two $\tau$-structures $(\mathfrak{A},\mathfrak{B})$ is a sequence ${\mathcal{H}}_{0},\mathcal{H}_{1},\dots,{\mathcal{H}}_{q}$ of non-empty sets of partial homomorphisms from $\mathfrak{A}$ to $\mathfrak{B}$ such that
@@
To prove the connection between GQ+ and the $q$ nested guarded $k$ pebble game we need an appropriate notion of canonical Boolean $q$ nested guarded queries. In analogy to Section 4, it suffices to present this connection for complements of CSPs of $\omega$-categorical structures $\mathfrak{B}$, by Corollary 3.19.
@@
Let $\mathfrak{B}$ be an $\omega$-categorical structure with a finite relational signature $\tau$ and let $R\subseteq B^{m}$ be a relation with a primitive positive definition over $\mathfrak{B}$. The canonical $q$ nested guarded query $\Pi_{R}$ of width $(m,k)$ for $(\mathfrak{B},R)$ is defined by induction over $q$ as follows. For $q=1$, let $o^{1}=(o^{1}_{1},\dots,o^{1}_{m}),\dots,o^{s}=(o^{s}_{1},\dots,o^{s}_{m})\in B^{m}$ be representatives of all orbits of $m$-tuples of $\operatorname{Aut}(\mathfrak{B})$ that are not contained in $R$. For $i\in\{1,\dots,s\}$ let $\Pi_{o^{i}}$ be the canonical frontier-guarded Datalog program of the $\omega$-categorical structure $(\mathfrak{B};o^{i}_{1},\dots,o^{i}_{m})$. We transform $\Pi_{o^{i}}$ into an $m$-ary FCP by replacing literals of the form $o^{i}_{j}=x$ by $\lambda_{j}=x$. Let $\Pi$ be obtained by taking the conjunction of all the resulting FCPs over all $i\in\{1,\dots,s\}$ (which is again an FCP by Lemma 4.3).
For $q>1$, we suppose inductively that for all $m\geq 0$ and relations $S$ with a primitive positive definition over $\mathfrak{B}$ the canonical $q-1$ nested guarded query $\Xi_{S}$ of width $(m,k)$ for $(\mathfrak{B},S)$ is already defined. Replace in the canonical $1$ nested guarded query of width $(m,k)$ for $(\mathfrak{B},R)$ every occurrence of an EDB $S$ by $\Xi_{S}$. We define $\Pi_{R}$ to be the resulting $q$-nested query. The canonical $q$ nested guarded query $\Pi$ of width $(m,k)$ for $\mathfrak{B}$ is obtained from the canonical $q$ nested guarded query $\Pi$ of width $(m,k)$ for $(\mathfrak{B},\emptyset)$ by existentially quantifying all free variables.
Definition 6.25 ($q$ nested $m$ bounded $k$ variable logic).
Let $\tau$ be a relational signature and let $q,m,k\in{\mathbb{N}}$. Then $\tau$-sentences in $q$ nested $m$ bounded $k$ variable logic $L^{m,k}_{q}$ are built inductively from atomic $\tau$-formulas and the following operation: if $\phi_{1},\dots,\phi_{s}$ are formulas from $L^{m,k}_{q}$, then
@@
    Every $q$ nested guarded query of width $(m,k)$ which is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$ is false on $\mathfrak{A}$.
@@
    Duplicator has a winning strategy for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{B})$.
@@
For the implication from 1. to 2., it suffices to prove that the canonical $q$-nested guarded query $\Pi$ of width $(m,k)$ is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$. Let $\mathfrak{A}$ be a finite $\tau$-structure. Let $R\subseteq B^{m}$ be primitively positively definable over $\mathfrak{B}$. First observe that if $(s_{1},\dots,s_{m})$ and $(s_{1}^{\prime},\dots,s_{m}^{\prime})$ have the same orbit in $\operatorname{Aut}(\mathfrak{B})$, then $(\mathfrak{B},s_{1},\dots,s_{m})$ and $(\mathfrak{B},s_{1}^{\prime},\dots,s_{m}^{\prime})$ have the same canonical frontier-guarded Datalog program. Hence, if $q=1$ then $\Pi(t_{1},\dots,t_{m})$ holds in $\mathfrak{A}$ if and only if for every tuple $(s_{1},\dots,s_{m})\in B^{m}\setminus R$ the canonical frontier-guarded Datalog program for $(\mathfrak{B},s_{1},\dots,s_{m})$ derives ${\sf goal}$. This in turn means that there exists no homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$ which maps $(t_{1},\dots,t_{m})$ to $(s_{1},\dots,s_{m})$. By induction over the nesting depth $q$ it follows that if $\mathfrak{A}$ satisfies the canonical $q$ nested guarded query of with $(m,k)$ for $\mathfrak{B}$ then there is no homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$, and hence $\Pi$ is sound for $\operatorname{CSP}(\mathfrak{B})$.
@@
For the implication from 2. to 3., let $\mathfrak{B}^{\prime}$ be the expansion of $\mathfrak{B}$ by all primitive positive definable relations of arity at most $m$; since $\mathfrak{B}$ is $\omega$-categorical, the structure $\mathfrak{B}^{\prime}$ still has a finite relational signature $\tau^{\prime}$. We compute a winning strategy ${\mathcal{H}}_{0},{\mathcal{H}}_{1},\dots,{\mathcal{H}}_{q}$ for Duplicator for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{B})$ using the canonical $q$-nested guarded query of width $(m,k)$ for $(\mathfrak{B}^{\prime},R)$ for every relation $R$ of arity $m$ which has a primitive positive definition in $\mathfrak{B}$. Define $\mathfrak{A}^{\prime}_{0}$ be the expansion of $\mathfrak{A}$ with the same signature as $\mathfrak{B}^{\prime}$ that contains for every $m$-ary relation symbol $R\in\tau^{\prime}\setminus\tau$ the empty relation, i.e., $R^{\mathfrak{A}^{\prime}_{0}}=\emptyset$. For $i\in\{0,\dots,q-1\}$, suppose that $\mathfrak{A}_{0}^{\prime},\dots,\mathfrak{A}_{i}^{\prime}$ have already been defined. Let $\mathfrak{A}^{\prime}_{i+1}$ be the expansion of $\mathfrak{A}$ with the same signature as $\mathfrak{B}^{\prime}$ that contains for every $m$-ary relation symbol $R\in\tau^{\prime}\setminus\tau$ the relation computed by the canonical $i$-nested guarded query of width $(m,k)$ for $(\mathfrak{B}^{\prime},R)$ on $\mathfrak{A}^{\prime}_{i-1}$. For $i\in\{0,\dots,q\}$, let ${\mathcal{H}}_{i}$ be the set of all partial homomorphisms $f$ from $\mathfrak{A}_{i}^{\prime}$ to $\mathfrak{B}^{\prime}$ with domain of size at most $k$. Note that since the canonical $q$ nested guarded query of width $(m,k)$ for $\mathfrak{B}$ is false on $\mathfrak{A}$, the relation symbol $F$ of arity 00 which denotes the empty relation over $\mathfrak{B}$ (which is primitively positively definable) denotes the empty relation over $\mathfrak{A}^{\prime}_{q}$ as well, and hence ${\mathcal{H}}_{0},{\mathcal{H}}_{1},\dots,{\mathcal{H}}_{q}$ are non-empty. Clearly, these sets are closed under restriction.
@@
We show item 3. by induction over $i\in\{1,\dots,q\}$. Let $h\in{\mathcal{H}}_{i}$ be such that $\operatorname{dom}(h)$ has size at most $m$ and let $S\subseteq A$ be a superset of $\operatorname{dom}(h)=\{a_{1},\dots,a_{j}\}$ of size $j\leq k$. We have to show that ${\mathcal{H}}_{i-1}$ contains an extension of $h$ with domain $S$. Let $\phi$ be the primitive positive formula obtained from the canonical query of the substructure $\mathfrak{A}_{S}$ of $\mathfrak{A}^{\prime}_{i}$ induced on $S$ by existentially quantifying all variables except the ones in the domain of $h$. Then $\phi$ is equivalent to a primitive positive formula over the signature of $\mathfrak{B}$; let $R$ be the relation symbol of $\mathfrak{B}^{\prime}$ for the relation defined by $\phi$ in $\mathfrak{B}$. For any $t\in B^{j}\setminus R^{\mathfrak{B}^{\prime}}$ the canonical $q$-nested query derives ${\sf goal}$ since the canonical query of $\mathfrak{A}_{S}$ together with $a_{1}=t_{1},\dots,a_{j}=t_{j}$ is unsatisfiable in $(\mathfrak{B},t_{1},\dots,t_{j})$. Hence, $a\in R^{\mathfrak{A}^{\prime}}$. In this case, $h(a)\in R^{\mathfrak{B}^{\prime}}$ since $h$ is a partial homomorphism on the entries of $a$. Therefore, $h(a)$ satisfies $\phi$ and the witnesses for the existentially quantified variables of $\phi$ provide an extension of $h$ to a partial homomorphism from $\mathfrak{A}^{\prime}_{i+1}$ to $\mathfrak{B}$ which is defined on all of $S$.
3. implies 4.: We show by induction over the syntactic structure of $L^{m,k}_{q}$ formulas that if $\phi(v_{1},\dots,v_{m})$ is an $L^{\ell,k}_{q}$ formula, then for all $h\in{\mathcal{H}}_{q}$ and all elements $a_{1},\dots,a_{m}$ from the domain of $h$, if $\mathfrak{A}$ satisfies $\phi(a_{1},\dots,a_{m})$, then $\mathfrak{B}$ satisfies $\phi(h(a_{1}),\dots,h(a_{m}))$. For $m=0$ this implies 4. The base case of the induction is obvious because atomic formulas are preserved by homomorphisms. Now suppose that $\phi$ is of the form $\exists\bar{y}(\phi_{1}\wedge\cdots\wedge\phi_{s})$ where $\phi_{1},\dots,\phi_{s}$ are formulas from $L^{m,k}_{q}$. Let $\bar{c}$ be a tuple providing witnesses for the variables $\bar{y}$ that shows that $\phi(a_{1},\dots,a_{m})$ holds in $\mathfrak{A}$.
@@
For each IDB $R$ of $\Pi$ of arity $k$ and every $k$-tuple $\bar{a}$ such that $\Pi$ derives $R(\bar{a})$ on $\mathfrak{A}$ we define an $L^{m,k}_{q}$-formula $\psi(x_{1},\dots,x_{k})$ which holds on $\bar{a}$ in $\mathfrak{A}$ and such that $\Pi$ derives $R(\bar{x})$ on the canonical database of $\psi(\bar{x})$ (after transforming it into prenex normal form) by induction over the evaluation of $\Pi$ on $\mathfrak{A}$. Let $\theta(x_{1},\dots,x_{k},\bar{y})$ be the body of the rule of $\Pi$ that derived $R(\bar{a})$ on $(\bar{a},\bar{b})$. For every atomic formula $S(\bar{z})$ of $\theta$ we may assume by inductive assumption that there exists an $L^{m,k}_{q}$-formula $\chi(\bar{z})$ which holds on the respective entries $\bar{c}$ of $(\bar{a},\bar{b})$ in $\mathfrak{A}$ and such that $\Pi$ derives $S(\bar{c})$ on the canonical database of $\chi(\bar{z})$. Then we define $\psi$ to be the conjunction of all these formulas where all variables $\bar{y}$ are existentially quantified. This formula clearly holds on $\bar{a}$ in $\mathfrak{A}$, and $\Pi$ derives $R^{(}\bar{x})$ on the canonical database of $\psi$. By the guard assumption, the rule of $\Pi$ that derives $R(\bar{a})$ must contain an atomic formula whose variables contain $x_{1},\dots,x_{k}$, so the formula $\psi$ is indeed an $L^{m,k}_{q}$-formula. Then the formula that we obtain for the IDB $R$ of $\Pi$ that denotes $B^{k}$ in $\mathfrak{B}^{\prime}$ by existentially quantifying all variables is true on $\mathfrak{A}$ but false on $\mathfrak{B}$.

### v3

@@ 5.4 Nested Queries
The expressive power of Datalog fragments can sometimes be increased by considering nesting, which is a concept introduced in [43] and further studied in [21]. The idea is that in nested queries we allow the use of other nested queries as if they were atomic formulas. Depending on the query language we start with, nesting may or may not increase the expressive power. Nested Datalog programs, for example, can be rewritten into equivalent Datalog programs without nesting. Similarly, MSO and GSO are closed under nesting. On the other hand, nested monadic queries (Nemodeq) are strictly more expressive than monadic queries [43], and, similarly, nested guarded queries (GQ+) are strictly more expressive than guarded queries, as we will see below (Proposition 5.31). We mention that the layered tree programs studied in [24] are in fact the queries obtained from linear monadic Datalog programs with at most one EDB per rule via nesting.
@@
Let $q,m\in{\mathbb{N}}$ and let $\tau$ be a finite relational signature. A $q$-nested (monadic, guarded) flag-and-check $\tau$-program of arity $m$ is defined inductively as follows. A $1$-nested (monadic, guarded) $\tau$-FCP of arity $m$ is the same as a (monadic, guarded) $\tau$-FCP of arity $m$ as defined in Definition 5.8. A $q+1$-nested (monadic, guarded) $\tau$-FCP of arity $m$ is a (monadic, guarded) $\tau$-FCP $\Pi$ of arity $m$ that may use in rule bodies $q$-nested (monadic, guarded) $\tau$-FCPs in addition to the symbols of $\tau$. These queries are called the subqueries of $\Pi$ at nesting level $q$; the subqueries of $\Pi$ at nesting level $j$ are inductively defined as the subqueries of the subqueries at nesting level $j+1$. In the case of guarded $q+1$-nested FCPs, the $q$-nested FCPs may not serve as guards. An FCP has width $m$ if all its subqueries have arity at most $m$.
@@
Note that all the classes of queries defined in Definition 5.16 are closed under disjunction by appropriately taking the union of the rule sets, just as in the proof of Lemma 4.3 for Datalog. $\triangle$
The class of nested guarded queries is denoted by GQ+, and the class of nested monadically defined queries by Nemodeq. It is known that every nested monadically defined query is equivalent to a Datalog program [43]; the same is even true for every nested guarded query [21, Figure 1]. The following fact is implicit in [21].
@@
  • replacing each atomic formula $R(\bar{y})$ by a new atomic formula $R^{\prime}(x_{1},\dots,x_{m},\bar{y})$ where $R^{\prime}$ is a new IDB.
For $q>1$, the translation is defined recursively: let $\Pi_{1},\dots,\Pi_{s}$ be the subqueries of $\Pi$ at nesting level q1q-1. Let $\Pi^{\prime}_{1},\dots,\Pi^{\prime}_{s}$ be the Datalog programs that can be associated to these FCPs by the inductive assumption; we may assume that the IDBs for all these programs are pairwise disjoint. Then $\Pi^{\prime}$ is defined to be the union over the rules of $\Pi_{1},\dots,\Pi_{s}$ together with the rules of $\Pi$ where we replace each subquery $\Pi_{i}$ at nesting level $q$ by $P_{\Pi_{i}}$.
@@
Example 5.11 can be expressed in Nemodeq [43], and hence is in Datalog and in MSO. $\triangle$
@@ 5.5 The Nested Guarded Game
In this section we present a modification of the existential guarded $k$-pebble game that allows us to capture the expressive power of GQ and GQ+. In particular, we will prove that the query in Example 5.11 is not in GQ. We call this game the nested guarded game; it will also be used in the next section to prove that there are problems in the intersection of Datalog and GSO that cannot be expressed in GQ+.
Again we start with an informal description of the game. Let $\tau$ be a finite relational signature. The game has three parameters, $q$, $m$, and $k$. There are two players, Spoiler and Duplicator, that play on a pair of $\tau$-structures $(\mathfrak{A},\mathfrak{B})$. Each player has $k$ labelled pebbles, out of which $m$ pebbles are blue, the others are red. Spoiler places the blue pebbles on $\mathfrak{A}$, Duplicator answers by placing her blue pebbles on elements of $\mathfrak{B}$. Then the two players play the existential guarded (km)(k-m)-pebble game with the red pebbles, with the difference that Duplicator loses if the map between all the pebbled vertices (blue and red alike) is not a partial homomorphism. At most $q$ times, Spoiler can relocate the red pebbles (without the guard restriction). We present a formal definition of winning strategies for Duplicator in the nested guarded game; all statements about the game only involve such winning strategies for Duplicator.
@@
Let $k,q,m\in{\mathbb{N}}$. A winning strategy for Duplicator for the $q$-nested width $m$ guarded $k$-pebble game on two $\tau$-structures $(\mathfrak{A},\mathfrak{B})$ is a sequence ${\mathcal{H}}_{0},\mathcal{H}_{1},\dots,{\mathcal{H}}_{q}$ of non-empty sets of partial homomorphisms from $\mathfrak{A}$ to $\mathfrak{B}$ such that
@@
To prove the connection between GQ+ and the $q$-nested guarded $k$-pebble game we need an appropriate notion of canonical Boolean $q$-nested guarded queries.
Remark 5.24.
We present the game characterisation of GQ+ only for such GQ+ sentences whose sets of finite models are complements of CSPs of $\omega$-categorical structures $\mathfrak{B}$; however, in analogy to Section 4, one can define a game that characterises the general case, using Corollary 3.24. Since this is notationally and technically heavier, but does not add new ideas or new applications in the following, we only present the special case. $\triangle$
@@
Let $\mathfrak{B}$ be an $\omega$-categorical structure with a finite relational signature $\tau$ and let $R\subseteq B^{m}$ be a relation with a primitive positive definition over $\mathfrak{B}$. The canonical $q$-nested guarded query $\Pi_{R}$ of width $(m,k)$ for $(\mathfrak{B},R)$ is defined by induction over $q$ as follows. For $q=1$, let $o^{1}=(o^{1}_{1},\dots,o^{1}_{m}),\dots,o^{s}=(o^{s}_{1},\dots,o^{s}_{m})\in B^{m}$ be representatives of all orbits of $m$-tuples of $\operatorname{Aut}(\mathfrak{B})$ that are not contained in $R$. For $i\in\{1,\dots,s\}$ let $\Pi_{o^{i}}$ be the canonical frontier-guarded Datalog program of width $k$ of the $\omega$-categorical structure $(\mathfrak{B};o^{i}_{1},\dots,o^{i}_{m})$. [^8] Observe that if $(s_{1},\dots,s_{m})$ and $(s_{1}^{\prime},\dots,s_{m}^{\prime})$ have the same orbit in $\operatorname{Aut}(\mathfrak{B})$, then $(\mathfrak{B},s_{1},\dots,s_{m})$ and $(\mathfrak{B},s_{1}^{\prime},\dots,s_{m}^{\prime})$ have the same canonical frontier-guarded Datalog program. We transform $\Pi_{o^{i}}$ into an $m$-ary FCP by replacing literals of the form $o^{i}_{j}=x$ by $\lambda_{j}=x$. Let $\Pi_{R}$ be obtained by taking the conjunction of all the resulting FCPs over all $i\in\{1,\dots,s\}$ (which is again an FCP; this can be shown analogously to Lemma 4.3).
For $q>1$, we suppose inductively that for all $m\geq 0$ and relations $S$ with a primitive positive definition over $\mathfrak{B}$ the canonical q1q-1 nested guarded query $\Xi_{S}$ of width $(m,k)$ for $(\mathfrak{B},S)$ is already defined. Replace in the canonical $1$ nested guarded query of width $(m,k)$ for $(\mathfrak{B},R)$ every occurrence of an EDB $S$ by $\Xi_{S}$. We define $\Pi_{R}$ to be the resulting $q$-nested query.
The canonical $q$ nested guarded query $\Pi$ of width $(m,k)$ for $\mathfrak{B}$ is obtained from the canonical $q$ nested guarded query $\Pi_{R}$ of width $(m,k)$ for $(\mathfrak{B},R)$, where $R$ is the empty relation of arity 0, by existentially quantifying all free variables.
Definition 5.26 ($q$-nested $m$-bounded $k$-variable logic).
Let $\tau$ be a relational signature and let $q,m,k\in{\mathbb{N}}$. Then $\tau$-sentences in $q$-nested $m$-bounded $k$-variable logic $L^{m,k}_{q}$ are built inductively from atomic $\tau$-formulas and the following operation: if $\phi_{1},\dots,\phi_{s}$ are formulas from $L^{m,k}_{q}$, then
@@
The following result and its proof are inspired by Theorem 4.2 and its proof.
@@
    Every $q$-nested guarded query of width $(m,k)$ which is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$ is false on $\mathfrak{A}$.
@@
    Duplicator has a winning strategy for the $q$-nested width $m$ guarded $k$-pebble game on $(\mathfrak{A},\mathfrak{B})$.
@@
For the implication from 1. to 2., it suffices to prove that the canonical $q$-nested guarded query $\Pi$ of width $(m,k)$ is sound for the complement of $\operatorname{CSP}(\mathfrak{B})$. Let $\mathfrak{A}$ be a finite $\tau$-structure. Let $R\subseteq B^{m}$ be primitively positively definable over $\mathfrak{B}$. If $q=1$ then $\Pi(t_{1},\dots,t_{m})$ holds in $\mathfrak{A}$ if and only if for every tuple $(s_{1},\dots,s_{m})\in B^{m}\setminus R$ the canonical frontier-guarded Datalog program for $(\mathfrak{B},s_{1},\dots,s_{m})$ derives $\sf goal$}. This in turn means that there exists no homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$ which maps $(t_{1},\dots,t_{m})$ to $(s_{1},\dots,s_{m})$, because the canonical frontier-guarded Datalog program, which is a subset of the canonical Datalog program of width $k$, is sound by definition. By induction over the nesting depth $q$ it follows that if $\mathfrak{A}$ satisfies the canonical $q$ nested guarded query of with $(m,k)$ for $\mathfrak{B}$, then there is no homomorphism from $\mathfrak{A}$ to $\mathfrak{B}$, and hence $\Pi$ is sound for $\operatorname{CSP}(\mathfrak{B})$.
@@
For the implication from 2. to 3., let $\mathfrak{B}^{\prime}$ be the expansion of $\mathfrak{B}$ by all primitive positive definable relations of arity at most $m$; since $\mathfrak{B}$ is $\omega$-categorical, the structure $\mathfrak{B}^{\prime}$ still has a finite relational signature $\tau^{\prime}$. We compute a winning strategy ${\mathcal{H}}_{0},{\mathcal{H}}_{1},\dots,{\mathcal{H}}_{q}$ for Duplicator for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{B})$ using the canonical $q$-nested guarded query of width $(m,k)$ for $(\mathfrak{B}^{\prime},R)$ for every relation $R$ of arity $m$ which has a primitive positive definition in $\mathfrak{B}$. Define $\mathfrak{A}^{\prime}_{0}$ be the expansion of $\mathfrak{A}$ with the same signature as $\mathfrak{B}^{\prime}$ that contains for every $m$-ary relation symbol $R\in\tau^{\prime}\setminus\tau$ the empty relation, i.e., $R^{\mathfrak{A}^{\prime}_{0}}=\emptyset$. For $i\in\{0,\dots,q-1\}$, suppose that $\mathfrak{A}_{0}^{\prime},\dots,\mathfrak{A}_{i}^{\prime}$ have already been defined. Let $\mathfrak{A}^{\prime}_{i+1}$ be the expansion of $\mathfrak{A}$ with the same signature as $\mathfrak{B}^{\prime}$ that contains for every $m$-ary relation symbol $R\in\tau^{\prime}\setminus\tau$ the relation computed by the canonical $i$-nested guarded query of width $(m,k)$ for $(\mathfrak{B}^{\prime},R)$ on $\mathfrak{A}^{\prime}_{i-1}$. For $i\in\{0,\dots,q\}$, let ${\mathcal{H}}_{i}$ be the set of all partial homomorphisms $f$ from $\mathfrak{A}_{i}^{\prime}$ to $\mathfrak{B}^{\prime}$ with domain of size at most $k$. Note that since the canonical $q$ nested guarded query of width $(m,k)$ for $\mathfrak{B}$ is false on $\mathfrak{A}$, the relation symbol $F$ of arity 0 which denotes the empty relation over $\mathfrak{B}$ (which is primitively positively definable) denotes the empty relation over $\mathfrak{A}^{\prime}_{q}$ as well, and hence ${\mathcal{H}}_{0},{\mathcal{H}}_{1},\dots,{\mathcal{H}}_{q}$ are non-empty. Clearly, these sets are closed under restriction.
@@
We show item 3. by induction over $i\in\{1,\dots,q\}$. Let $h\in{\mathcal{H}}_{i}$ be such that $\operatorname{dom}(h)$ has size at most $m$ and let $S\subseteq A$ be a superset of $\operatorname{dom}(h)=\{a_{1},\dots,a_{j}\}$ of size $j\leq k$. We have to show that ${\mathcal{H}}_{i-1}$ contains an extension of $h$ with domain $S$. Let $\phi$ be the primitive positive formula obtained from the canonical query of the substructure $\mathfrak{A}_{S}$ of $\mathfrak{A}^{\prime}_{i}$ induced on $S$ by existentially quantifying all variables except the ones in the domain of $h$. Then $\phi$ is equivalent to a primitive positive formula over the signature of $\mathfrak{B}$; let $R$ be the relation symbol of $\mathfrak{B}^{\prime}$ for the relation defined by $\phi$ in $\mathfrak{B}$. For any $(t_{1},\dots,t_{j})\in B^{j}\setminus R^{\mathfrak{B}^{\prime}}$ the canonical $q$-nested query derives $\sf goal$} since the canonical query of $\mathfrak{A}_{S}$ together with $a_{1}=t_{1},\dots,a_{j}=t_{j}$ is unsatisfiable in $(\mathfrak{B},t_{1},\dots,t_{j})$. Hence, $(a_{1},\dots,a_{j})\in R^{\mathfrak{A}^{\prime}}$. In this case, $(h(a_{1}),\dots,h(a_{j}))\in R^{\mathfrak{B}^{\prime}}$ since $h$ is a partial homomorphism on $\{a_{1},\dots,a_{j}\}$. Therefore, $(h(a_{1}),\dots,h(a_{j}))$ satisfies $\phi$ and the witnesses for the existentially quantified variables of $\phi$ provide an extension of $h$ to a partial homomorphism from $\mathfrak{A}^{\prime}_{i+1}$ to $\mathfrak{B}$ which is defined on all of $S$.
3. implies 4.: We show by induction over the syntactic structure of $L^{m,k}_{q}$ formulas that if $\phi(v_{1},\dots,v_{m})$ is an $L^{m,k}_{q}$ formula, then for all $h\in{\mathcal{H}}_{q}$ and all elements $a_{1},\dots,a_{m}$ from the domain of $h$, if $\mathfrak{A}$ satisfies $\phi(a_{1},\dots,a_{m})$, then $\mathfrak{B}$ satisfies $\phi(h(a_{1}),\dots,h(a_{m}))$. For $m=0$ this implies 4. The base case of the induction is obvious because atomic formulas are preserved by homomorphisms. Now suppose that $\phi$ is of the form $\exists\bar{y}(\phi_{1}\wedge\cdots\wedge\phi_{s})$ where $\phi_{1},\dots,\phi_{s}$ are formulas from $L^{m,k}_{q}$. Let $\bar{c}$ be a tuple providing witnesses for the variables $\bar{y}$ that shows that $\phi(a_{1},\dots,a_{m})$ holds in $\mathfrak{A}$.
@@
For each IDB $R$ of $\Pi$ of arity $k$ and every $k$-tuple $\bar{a}$ such that $\Pi$ derives $R(\bar{a})$ on $\mathfrak{A}$ we define an $L^{m,k}_{q}$-formula $\psi(x_{1},\dots,x_{k})$ which holds on $\bar{a}$ in $\mathfrak{A}$ and such that $\Pi$ derives $R(\bar{x})$ on the canonical database of $\psi(\bar{x})$ (after transforming it into prenex normal form) by induction over the evaluation of $\Pi$ on $\mathfrak{A}$. Let $\theta(x_{1},\dots,x_{k},\bar{y})$ be the body of the rule of $\Pi$ that derived $R(\bar{a})$ on $(\bar{a},\bar{b})$. For every atomic formula $S(\bar{z})$ of $\theta$ we may assume by inductive assumption that there exists an $L^{m,k}_{q}$-formula $\chi(\bar{z})$ which holds on the respective entries $\bar{c}$ of $(\bar{a},\bar{b})$ in $\mathfrak{A}$ and such that $\Pi$ derives $S(\bar{c})$ on the canonical database of $\chi(\bar{z})$. Then we define $\psi$ to be the conjunction of all these formulas where all variables $\bar{y}$ are existentially quantified. This formula clearly holds on $\bar{a}$ in $\mathfrak{A}$, and $\Pi$ derives $R^{(}\bar{x})$ on the canonical database of $\psi$. The rule of $\Pi$ that derives $R(\bar{a})$ must be guarded, and hence contains an atomic formula whose variables contain $x_{1},\dots,x_{k}$, so the formula $\psi$ is indeed an $L^{m,k}_{q}$-formula. Then the formula that we obtain for the IDB $R$ of $\Pi$ that denotes $B^{k}$ in $\mathfrak{B}^{\prime}$ by existentially quantifying all variables is true on $\mathfrak{A}$ but false on $\mathfrak{B}$.